OpenStudy (anonymous):

Step by step solve plz: sqrt of -32 (I know Im stupid,dont rub it in)

OpenStudy (anonymous):

32 = 2*2*2*2*2 -32 = -2*-2*-2*-2*-2 So, sqrt -32 = -4 sqrt 2 \[-4\sqrt{2}\]

OpenStudy (anonymous):

................even tho I dont get the process,its an answer,so thanx..

OpenStudy (anonymous):

Welcome.

OpenStudy (anonymous):

I'm sorry, I made a mistake, its 4sqrt2.